OpenStudy (anonymous):

Simplify (6y^2 -6/ 8y^2 +8y)/(3y -3 / 4y^2 + 4y)

OpenStudy (anonymous):

There you can use this: \[\frac{\frac{A}{B}}{\frac{C}{D}}=\frac{AD}{BC}\]

OpenStudy (anonymous):

Where: A=6y^2 -6 B=8y^2 +8y C=3y -3 D=4y^2 + 4y What do we get?

OpenStudy (anonymous):

a big number lol

OpenStudy (anonymous):

:D Best answer for you!

OpenStudy (anonymous):

no!!!

OpenStudy (anonymous):

Could you continue? What did you get?

OpenStudy (anonymous):

21y+32/28-3y i think

OpenStudy (anonymous):

\[ \frac{(6y^2 -6)(4y^2 + 4y)}{(8y^2 +8y)(3y -3)}=\frac{(6y^2-6)}{2(3y-3)}\]

OpenStudy (anonymous):

I got that.

OpenStudy (anonymous):

so i got it wrong

OpenStudy (anonymous):

dont you have to solve that now

OpenStudy (anonymous):

Yes, we can do this: \[\frac{6y^2-6}{2(3y-3)}=\frac{6(y-1)}{6y-6}=\frac{y^2-1}{y-1}\] And we can do one more step, which one?

OpenStudy (anonymous):

take out the 1s

OpenStudy (anonymous):

and y^2-y^1

OpenStudy (anonymous):

We can use this: \[y^2-1=(y+1)(y-1)\]

OpenStudy (anonymous):

this math is so hard

OpenStudy (anonymous):

Pluging that on the numerator., we get this: \[\frac{(y+1)(y-1)}{y-1}\] So, the final answer is?

OpenStudy (anonymous):

y-1

OpenStudy (anonymous):

Sure?

OpenStudy (anonymous):

we can cancell out y-1

OpenStudy (anonymous):

@brendas The final answer is y+1

OpenStudy (anonymous):

i knew that thanks

OpenStudy (anonymous):

OK, awsome! i hope I have helped you tonight!

OpenStudy (anonymous):

PD: Thanks for the "best response"!! :D